OpenStudy (anonymous):

Approximate to the nearest tenth the real zeros of f(x) = –5x3 + 9x2 + 12x + 2. answers: The zeros are approximately 0.2, 0.7, and 2.7. The zeros are approximately –2.7, –1.7, and 0.2. The zeros are approximately –0.2, –0.7, and 2.7. The zeros are approximately –0.5, –0.7, and 1.7.

OpenStudy (shubhamsrg):

try sum of roots thing i guess..

OpenStudy (anonymous):

\[-5 x^3 + 9 x^2 +12 x+ 2=(-5 x-1) \left(x^2-2 x-2\right) \]

OpenStudy (anonymous):

what is that @ robtobey thats not the one of th answers

OpenStudy (anonymous):

or is it?

OpenStudy (anonymous):

\[\left\{-\frac{1}{5},1-\sqrt{3},1+\sqrt{3}\right\}\to \{-0.2,-0.732051,2.73205\} \]Sorry, There was a sign mistake for the 1/5

OpenStudy (anonymous):

robtobey may you please explain how you guess roots ? (i heard of a method of some divisors of one of the numbers in the expression ?)

OpenStudy (anonymous):

i mean roots for the factorization

OpenStudy (anonymous):

Used Mathematica's Factor function. Refer to the attachment. Mathematica leaves the terms in reverse order, still valid though.

OpenStudy (perl):

this is a cubic so in theory it can be solved exactly. why dont you use your graphing calculator,

OpenStudy (anonymous):

dont know how to plug it in dude

OpenStudy (anonymous):

thanks dude

OpenStudy (perl):

:)

OpenStudy (anonymous):

For the general non factorable cubic Mathematica provides the NSolve function. Refer to the attachment.

OpenStudy (anonymous):

ok so you didnt do it by hand

OpenStudy (anonymous):

No. I am not a student in school. http://www.ted.com/talks/conrad_wolfram_teaching_kids_real_math_with_computers.html